OpenStudy (anonymous):

What percent of 26 is 4? I will give medal (:

OpenStudy (anonymous):

4% of 26 is 1.04

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yes, it is 1.04

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Its not a choice..its either 0.26 = n × 4 n = 0.26 × 4 4 = n × 26 26 = 4n :\

OpenStudy (anonymous):

i think its n=0.26x4.. i think